• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Carlos

    Log entry time 08:02:17 on January24,2009

    Entry number 257553

    keyword=Owl shift summary


    Pretty smooth shift, only disturbed by some BigBite HV trips (which disappeared after adjusting beam position to the center of the target cell) and the alarms of Q1 (still noisy as yesterday). Happex roc28 crashed a few times, too. We took production data for most of the shift. A few events are detailed below.

    1.20 some bigbite hv trips

    2.00 MCC called. Bringing Hall C up, beam will be off for us for about 5min.

    2.15 Beam back. Resuming production.

    3.00 BB HV trip

    3.45 Jin called MCC to change beam position to reduce BB trips. New values: BPMA (-1.3,2.5), BPMB(-0.8,2.5).